romana is thinking about investing her savings in a mutual fund.. she found data in published fund report on various fund categories and estimated beta specific to its fund category for three funds she is considering as potential investment opportunities.compute the expected return for each of the three mutual funds listed below. using the capital asset pricing model and 2.85% as the rate of return on a risk free asset3:34 PMFund name a , beta 0.69 , expected return 13.503:35 PMfund name b, beta 1.19, expected return 6.143:36 PMfund name c , beta 2.06 , expeted return 12.30
Answer #1

As per CAPM

Expected return = Rf+ Beta*(Rm- Rf)

For fund 3:34 PM

Expected return = 2.85%+ 0.69*(13.5%-2.85%) =10.2%

3:35 PMfund

Expected return = 2.85%+ 1.19*(6.14%-2.85%) = 6.77%

3:36 Fund

Expected return = 2.85%+ 2.06*(12.3%-2.85%) = 22.32%